OpenStudy (anonymous):

a. Determine whether the ratios 10/20 and 30/60 can form a proportion by finding a common multiplier. b. Show that the ratios in part (a) are equal by writing them in simplest form. Please explain how you got the answers to these questions.

OpenStudy (anonymous):

You know how to find common multipliers for just two numbers?

OpenStudy (anonymous):

b) \[\frac{ 10 }{ 20 }=\frac{ 30 }{60 }=>10*60=20*30\] which is true